# AutoFT: Robust Fine-Tuning by Optimizing Hyperparameters on OOD Data

### 18 Jan 2024 v1

@@ 4.2 Hyperparameter Optimization Loop
Formulation. We formalize our hyperparameter optimization as follows. Given a hyperparameter space $\Phi$ for a fine-tuning algorithm LearnAlg, we solve the optimization problem in 1 to find hyperparameters $\phi^{*}$ such that the fine-tuned model $f^{\prime}=\text{LearnAlg}(\phi^{*},D_{\text{tr}})$ performs well on the OOD validation set $D_{\text{val}}$. Practically, this involves fine-tuning the same foundation model multiple times, each time with different hyperparameters $\phi^{1},\phi^{2},\ldots$ to obtain different final fine-tuned models $f^{1}=\text{LearnAlg}(\phi^{1},D_{\text{tr}}),f^{2}=\cdots$. We then judge the quality of each set of hyperparameters $\phi^{i}$ by the performance of $f^{i}$ on the OOD validation set. The hyperparameter optimization algorithm aggregates the information from these different fine-tuning runs and estimates the best hyperparameters $\phi^{*}$. At test time, we fine-tune the initial foundation model with hyperparameters $\phi^{*}$, and evaluate its performance on novel OOD distributions to assess its generalization capabilities. We summarize this procedure in Algorithm 1.
@@
Hyperparameter space. We consider a larger hyperparameter space $\Phi$ than what is typically considered. This is because for robust fine-tuning, the algorithm must selectively ignore some aspects of the training distribution in order to better generalize to novel OOD distributions. To enable this more nuanced behavior, we increase the expressivity of the hyperparameter optimization process by expanding the hyperparameter space $\Phi$. The main novelty of our hyperparameters space is that we allow the hyperparameter space $\Phi$ to express the fine-tuning objective itself by learning to weight between different loss functions and regularizers. For example, the cross-entropy and hinge losses similarly guide the model’s predictions towards the correct label, but with a different amount of penalty for severely incorrect samples. As another example, contrastive loss on image-text pairs [72, 30] also guides the model’s predictions towards the correct label, but in a way that incorporates language embeddings which is entirely different from how the cross-entropy and hinge losses operate. Similarly, different losses for reducing overconfidence or constraining model parameters towards their initial values have different effects on the model’s adaptation process.
@@
As hyperparameters, we consider weight coefficients for nine different loss functions and regularizers: cross-entropy loss, hinge loss, image-text contrastive loss, entropy, confidence minimization, L1 norm, L2 norm, L1 and L2 distance to pretrained model parameters. We select these losses to address various aspects of model learning and generalization, such as classification accuracy, decision confidence, and prevention of overfitting, ensuring a nuanced and effective model adaptation. We denote these loss weight coefficients as $W=\{w_{1},w_{2},\dots,w_{9}\}$. Denoting the $i$-th loss function or regularizer as $\mathcal{L}_{i}$, the total loss $\mathcal{L}$ is the weighted sum $\mathcal{L}=\sum_{i=1}^{9}w_{i}\mathcal{L}_{i}$. To enable finer control over the speed and stability of adaptation during fine-tuning, we additionally learn hyperparameters for the fine-tuning optimizer: learning rate $\eta$ and weight decay $\delta$. For an even larger hyperparameter space, we optionally consider per-layer hyperparameters for the learning rate, weight decay, and parameter norm weights; we call this variant “per-layer” in our experiments in Appendix A. Our complete set of hyperparameters is thus $\phi=(W,\eta,\delta)$.
@@
Hyperparameter optimization algorithm. Finding effective hyperparameters for robust fine-tuning requires an optimizer that can efficiently search our large, high-dimensional hyperparameter space $\Phi$. We use the Tree-structured Parzen Estimator [12, TPE] as the hyperparameter optimization algorithm, as its design allows for efficient processing in high-dimensional spaces and yields effective results at the order of hundreds of evaluations. TPE creates a probabilistic model from previous hyperparameter evaluations, and differentiates between better and worse hyperparameter configurations by comparing their likelihood under two separate models. We use the TPE implementation of the optuna library [2] with no further modifications.
@@
As the prior distribution for TPE, we use an appropriately scaled distribution for each hyperparameter, operating in the log-space for the weight coefficients and the learning rate since the optimal values can vary over several orders of magnitude:
$w_{i}\sim\text{LogUniform}$ $(w_{\text{min}},w_{\text{max}}),$
$\eta\sim\text{LogUniform}(\eta_{\text{min}},\,\eta_{\text{max}}),\,$ $\delta\sim\text{Uniform}(0.0,1.0).$
@@
Computational cost. The hyperparameter tuning of AutoFT introduces minimal computational overhead, requiring at most $5\%$ additional compute compared to one standard fine-tuning run. Each evaluation of hyperparameters on the OOD validation set requires only a few inner loop gradient steps on the fine-tuning dataset. For example, on the iWildCam dataset, AutoFT uses only 500 hyperparameter evaluations, each requiring only 10 gradient steps on the fine-tuning data. With 1.3 million training examples and a batch size of 256, AutoFT requires only 5000 additional gradient steps, compared to approximately 100,000 gradient steps for final fine-tuning. AutoFT is therefore computationally inexpensive, and highly practical for even large fine-tuning problem settings.
4.3 Intuition and Didactic Experiment
To provide intuition for how hyperparameter optimization can lead to more effective utilization of a partially useful prior, we consider a simple experiment. We focus on synthetic datasets of vectors drawn from zero-mean Gaussian distributions with different variances. We assume that the prior unit variance distribution is helpful but that the data exhibits deviation from this prior along a few dimensions. A detailed summary of our data and training setup is in Sec. B.1. We consider “fine-tuning” as starting from the prior distribution. In this toy model, dimension-wise weight hyperparameters $w_{i}$ dictate the degree of influence of the prior versus the training data on each dimension, analogous to how the fine-grained hyperparameters in AutoFT balance the foundation model prior and training data along different aspects of fine-tuning. As in AutoFT, we use the TPE algorithm to optimize the weight hyperparameters $w_{i}$ to maximize the log-likelihood of the validation data $D_{\text{val}}$.
@@ 5 Experimental Setup
Distribution shifts. In our analysis, we focus on natural distribution shifts, defined by Taori et al. [83] as shifts arising from real-world variations, such as changes in lighting, geography, and image styles. This approach, in line with prior works on robust fine-tuning [72, 92, 50, 30], emphasizes shifts that are representative of real-world scenarios.
Our main results include four distribution shifts, including real-world distribution shifts arising in wildlife recognition and satellite imagery, and two CIFAR-10-derived distribution shifts. The WILDS-iWildCam dataset [9, 47, 79] presents distribution shifts arising from variations in camera trap locations, lighting conditions, and animal behaviors across various geographic regions. The WILDS-FMoW dataset [16, 47, 79] presents distribution shifts arising from changes in time, geographic locations, and land use. Finally, we evaluate on two CIFAR-10 derived distribution shifts: CIFAR-10 to CIFAR-10.1 and CIFAR-10.2, which involve subtle changes in image characteristics and composition.
@@
Foundation models and CLIP. We fine-tune pre-trained CLIP models  [72], including those provided in the open-clip repository [40]. We use the CLIP ViT-B/16 model from OpenAI as our default model, unless specified otherwise. Evaluation on CIFAR-10 uses the CLIP ViT-L/14 model from OpenAI, in line with [92]. Our SoTA results on WILDS-iWildCam use the CLIP ViT-L/14-336px model from OpenAI. We use text templates used in prior work [72, 92] to generate zero-shot final layer weights for all datasets.
Effective robustness and weight ensembling curves. We use the effective robustness framework by Taori et al. [83] to evaluate model robustness based on accuracy exceeding a baseline trained only on the reference distribution. Linearly interpolating the weights of a fine-tuned model and a pretrained model (WiSE-FT) has been shown to improve both ID and OOD performance [91]. Hence, we include weight ensembling as an additional point of comparison, and interpolate the weights of models fine-tuned by each method with 10 mixing coefficients $\alpha$.
@@
Baselines. We compare AutoFT against several methods for adapting pretrained models. We include two standard transfer learning methods that minimize cross-entropy loss: linear probing (LP) and full fine-tuning (FT). We also compare with recent works in robust fine-tuning: L2-SP [59], which fine-tunes with an L2 regularization term towards pretrained weights; LP-FT [50], which performs linear probing followed by full fine-tuning; Freeze-Embed [51], which freezes the embedding layer during fine-tuning; and FLYP [30], which fine-tunes with the CLIP pretraining loss – a contrastive loss between image embeddings and class-descriptive prompt embeddings. As we are interested in OOD performance, we also compare against OOD generalization methods including Group DRO [78], ABSGD [70], LISA [95], DFR [44], and Copy-Paste [27]. We additionally evaluate all methods with weight ensembling (WiSE-FT) [92], which is shown to improve OOD performance in an orthogonal way to other robust fine-tuning methods.
Training protocol. We closely follow the training details of Goyal et al. [30] and Wortsman et al. [92]. All methods fine-tune models with an AdamW optimizer, cosine learning rate scheduler, and a batch size of 512 for ImageNet and 256 for all other datasets. All baseline hyperparameters, such as learning rate, weight decay, and warmup length, are tuned through grid search. All methods, including AutoFT, perform early stopping based on in-distribution (ID) validation accuracy. We provide a comprehensive breakdown of the hyperparameter sweeps in the supplementary material. We emphasize that none of these methods, including AutoFT, observe any of the test OOD distributions during training. Finally, we report metrics averaged over 5 runs with 95% confidence intervals.
@@ 6 Results
Table 1: iWildCam results. AutoFT with weight ensembling attains state-of-the-art OOD performance on the WILDS-iWildCam benchmark with a ViT-L/14-336px backbone, surpassing all prior entries on the WILDS leaderboard [47].
@@
Table 2: FMoW results. AutoFT with weight ensembling attains state-of-the-art OOD performance on the WILDS-FMoW benchmark with a ViT-L/14-336px backbone, surpassing all prior entries on the WILDS leaderboard [47].
@@
Table 3: AutoFT outperforms all baselines both with and without ensembling. Without ensembling, AutoFT improves OOD performance by $1.3\%$ on ImageNet, $3.7\%$ on WILDS-iWildCam, and $6.1\%$ on WILDS-FMoW. These improvements are preserved with weight ensembling.
Figure 4: AutoFT outperforms existing methods, both with and without weight ensembling [92]. Here, we show the ID-OOD performance curves obtained by linearly interpolating the fine-tuned model weights with the zero-shot weights.
In this section, we present the main experimental findings for AutoFT. First, we show that AutoFT improves the performance of fine-tuned models on several large-scale, synthetic and natural distribution shifts, including ImageNet, WILDS-iWildCam, WILDS-FMoW, and CIFAR. Then, we present additional experiments in the low-data and transfer learning regimes. Finally, we investigate the transferability of hyperparameters learned by AutoFT across fine-tuning dataset and backbone, and the effect of the choice of OOD validation distribution for hyperparameter optimization. These findings highlight the effectiveness of AutoFT for robust fine-tuning in a variety of settings.
@@
Table 4: AutoFT shows superior performance in binary few-shot classification. AutoFT outperforms FLYP by $3.1\%$ and full fine-tuning by $4.3\%$ in 32-shot classification on PatchCamelyon.

### 07 Mar 2024 v2

@@ 4.2 Learning a Fine-Tuning Objective via Bi-Level Optimization
Formulation. We frame the problem of learning a fine-tuning objective as a bi-level optimization problem with the form 1, which we can solve using existing black-box hyperparameter optimizers. Practically, solving this bi-level optimization problem involves fine-tuning the same foundation model multiple times, each time with different “hyperparameters” $\phi^{1},\phi^{2},\ldots$ to obtain different final fine-tuned models $f^{1}=\text{LearnAlg}(\phi^{1},D_{\text{tr}}),f^{2}=\cdots$. We then evaluate the quality of each $\phi^{i}$ through the performance of $f^{i}$ on the OOD validation set; by aggregating the information from these different fine-tuning runs, we can fine good hyperparameters $\phi^{*}$. At test time, we fine-tune the initial foundation model with $\phi^{*}$ and evaluate its performance on novel OOD distributions to assess its generalization capabilities. We summarize this procedure in Algorithm 1.
@@
Parameterization of the fine-tuning objective. The main novelty of AutoFT is that we parameterize the fine-tuning objective, which allows outer-loop optimization to search over an expressive space of fine-tuning algorithms. This extra expressivity is important: for robust fine-tuning, the algorithm must selectively ignore some aspects of the training distribution in order to better generalize to novel OOD distributions. Specifically, we allow $\Phi$ to express the fine-tuning objective by learning weights for a pre-defined set of loss functions and regularizers. Note that the cross-entropy and hinge losses similarly guide the model’s predictions towards the correct label, but with a different amount of penalty for severely incorrect samples. As another example, the contrastive loss on image-text pairs (Radford et al., 2021b; Goyal et al., 2022a) also guides the model’s predictions towards the correct label, but in a way that incorporates language embeddings, which is entirely different from how the cross-entropy and hinge losses operate.
@@
Therefore, we consider weight coefficients for nine different loss functions and regularizers: cross-entropy loss, hinge loss, image-text contrastive loss, entropy, confidence minimization, L1 norm, L2 norm, L1 and L2 distance to pretrained model parameters. We select these losses to address various aspects of model learning and generalization, such as classification accuracy, decision confidence, and prevention of overfitting, ensuring a nuanced and effective model adaptation. We denote these loss weight coefficients as $W=\{w_{1},w_{2},\dots,w_{9}\}$. Denoting the $i$-th loss function or regularizer as $\mathcal{L}_{i}$, the total loss $\mathcal{L}$ is the weighted sum $\mathcal{L}=\sum_{i=1}^{9}w_{i}\mathcal{L}_{i}$. To enable finer control over the speed and stability of adaptation during fine-tuning, we additionally learn hyperparameters for the fine-tuning optimizer: learning rate $\eta$, weight decay $\delta$, and random seed $\sigma$. Our complete set of hyperparameters for outer-loop optimization is thus $\phi=(W,\eta,\delta,\sigma)$.
@@
Bi-level optimization algorithm. We need a black-box optimizer that can effectively search through the high-dimensional hyperparameter space $\Phi$. We use the Tree-structured Parzen Estimator (Bergstra et al., 2011, TPE) for this purpose, as its design allows for efficient processing in high-dimensional spaces and yields effective results at the order of hundreds of evaluations. TPE creates a probabilistic model from previous hyperparameter evaluations, and differentiates between better and worse hyperparameter configurations by comparing their likelihood under two separate models. We use the TPE implementation from the optuna library (Akiba et al., 2019). Experiments in Table 8 demonstrate that in the context of learning a fine-tuning objective, TPE outperforms other bi-level optimizers such as random search, quasi Monte Carlo, and Bayesian optimization.
@@
As the prior distribution for TPE, we use an appropriately scaled distribution for each hyperparameter, operating in log-space for the weight coefficients and learning rate since the optimal values can vary over several orders of magnitude:
$w_{i}\sim\text{LogUniform}(w_{\text{min}},w_{\text{max}}),\quad$ $\rho\sim[0,100]$
$\eta\sim\text{LogUniform}(\eta_{\text{min}},\eta_{\text{max}}),\quad$ $\delta\sim\text{Uniform}(0.0,1.0).$
@@
Intuition for OOD evaluation. Optimizing the objective with respect to performance on a small out-of-distribution (OOD) validation set $D_{\text{val}}$ can guide the model to learn more generalizable parameters during fine-tuning. We conduct a didactic toy experiment in Figure 6, which illustrates our key intuition: a small validation set reveals how the data distribution changes under natural shifts, informing how best to adapt a given foundation model. Our experiments in Section 6 confirm that AutoFT consistently improves generalization to novel OOD distributions, outperforming state-of-the-art methods.
Computational cost. The bi-level optimization of AutoFT introduces minimal computational overhead, requiring at most $5\%$ additional compute compared to one standard fine-tuning run on most benchmarks. Each evaluation of the objective and hyperparameters requires a small number of inner loop gradient steps on the fine-tuning dataset. AutoFT is therefore computationally inexpensive, and highly practical for even large fine-tuning problem settings. Hyper-hyperparameters are detailed in Section D.4.
@@ 5 Experimental Setup
Evaluation. We fine-tune pre-trained CLIP models from the open-clip repository (Radford et al., 2021b; Ilharco et al., 2021). We use the CLIP ViT-B/16 model from OpenAI as our default model, unless specified otherwise. Our SoTA results on WILDS-iWildCam and WILDS-FMoW use the CLIP ViT-L/14-336px model. We use text templates used in prior work (Radford et al., 2021b; Wortsman et al., 2022b) to generate zero-shot classification weights.
Following prior works on robust fine-tuning, we evaluate AutoFT on nine natural distribution shifts, few-shot learning, and transfer learning settings (Taori et al., 2020; Radford et al., 2021b; Wortsman et al., 2022b; Kumar et al., 2022a; Goyal et al., 2022a). These include five ImageNet distribution shifts (IN-V2, IN-R, IN-Sketch, ObjectNet, and IN-A) (Recht et al., 2019; Hendrycks et al., 2021a; Wang et al., 2019; Barbu et al., 2019; Hendrycks et al., 2021b), WILDS-iWildCam (Koh et al., 2021; Beery et al., 2021), WILDS-FMoW (Koh et al., 2021; Christie et al., 2018), CIFAR-10.1 (Recht et al., 2018; Torralba et al., 2008), and CIFAR-10.2 (Lu et al., 2020). We use up to $1000$ examples from ImageNet-C, OOD validation splits of WILDS-FMoW and WILDS-iWildCam, and CIFAR-10-C as the OOD validation sets. We also evaluate on few-shot classification (Figure 5 and Figure 7) and transfer learning (Table 10).
Dataset Steps Trials $N_{\textrm{val}}$
iWildCam 10 500 1000
FMoW 10 500 1000
CIFAR-10 10 100 100
Flowers 50 500 500
Cars 50 500 1000
ImageNet 100 500 1000
Table 1: AutoFT training settings for each dataset.
Effective robustness and weight ensembling curves. We primarily evaluate robustness to distribution shifts using average OOD performance. We also compare methods with weight ensembling of the zero-shot and fine-tuned models (Wortsman et al., 2022b), which enhances ID and OOD performance in a way orthogonal to fine-tuning. We interpolate model weights with ten mixing coefficients, $\alpha$, and report results using the coefficient that maximizes ID validation accuracy. Finally, we plot ID and OOD performance across these different interpolation coefficients in Figure 3 to visualize effective robustness – gains in accuracy beyond the zero-shot model, which corresponds to vertical distance between curves.
@@
Baselines. We compare AutoFT against several methods for adapting pretrained models. We include two standard transfer learning methods that minimize cross-entropy loss: linear probing (LP) and full fine-tuning (FT). We also compare with recent works in robust fine-tuning: L2-SP (Li et al., 2018b), which fine-tunes with an L2 regularization term towards pretrained weights; LP-FT (Kumar et al., 2022a), which performs linear probing followed by full fine-tuning; Freeze-Embed (Kumar et al., 2022b), which freezes the embedding layer during fine-tuning; and FLYP (Goyal et al., 2022a), which fine-tunes with the CLIP pretraining loss – a contrastive loss between image embeddings and class-descriptive prompt embeddings. As we are interested in OOD performance, we also compare against OOD generalization methods, including Group DRO (Sagawa et al., 2019), ABSGD (Qi et al., 2020), LISA (Yao et al., 2022), DFR (Kirichenko et al., 2022), and Copy-Paste (Gao et al., 2023). We additionally evaluate all methods with weight ensembling (WiSE-FT) (Wortsman et al., 2022b), which is shown to improve OOD performance in an orthogonal way to other robust fine-tuning methods.
Bi-level optimization details. We summarize the number of inner-loop gradient steps, outer-loop fine-tuning trials, and validation set size for each dataset in Table 1. All of these values were selected based on performance on an ID validation set. We provide additional details on our training protocol in Section D.4.
@@ 6 Results
Table 2: iWildCam SoTA results. AutoFT with weight ensembling attains state-of-the-art OOD performance on the WILDS-iWildCam benchmark with a ViT-L/14-336px backbone, surpassing all prior entries on the WILDS leaderboard (Koh et al., 2021).
In this section, we present the main experimental findings for AutoFT. We show that AutoFT improves OOD generalization on nine distribution shifts. We present additional results in the low-data and IID transfer learning regimes in Section 6.2 and Section 6.3. These findings highlight the effectiveness of AutoFT for robust fine-tuning in a variety of settings. Finally, we investigate several aspects of the AutoFT method, including transferability of the learned objective across datasets, the choice of validation dataset and hyperparameter optimization algorithm. Further experimental details are in Appendix D.
@@
Table 3: FMoW SoTA results. AutoFT with weight ensembling attains state-of-the-art OOD performance on the WILDS-FMoW benchmark with a ViT-L/14-336px backbone, surpassing all prior entries on the WILDS leaderboard (Koh et al., 2021).
Figure 3: AutoFT outperforms existing methods, both with and without weight ensembling (Wortsman et al., 2022b). Here, we show the ID-OOD performance curves obtained by linearly interpolating weights of the fine-tuned model weights with the zero-shot model.
@@
Table 4: AutoFT outperforms all baselines, both with and without ensembling. Without ensembling, AutoFT improves OOD performance by $1.3\%$ on ImageNet, $2.7\%$ on WILDS-iWildCam, and $1.0\%$ on WILDS-FMoW. These improvements are preserved with weight ensembling.
Method ID OOD
FT 48.1 35.0
AutoFT, transferred objective 49.9 35.5
AutoFT 51.0 38.3
Table 5: Learning a fine-tuning objective on the task of interest (AutoFT) outperforms fine-tuning with an objective learned on a different task (AutoFT, transferred objective). This suggests that the learned objective is not universal, but specific to the task on which it is optimized.
Method Validation Set OOD Test
FT CIFAR-10 93.6 (0.2)
AutoFT CIFAR-10 95.1 (0.2)
CIFAR-10-C 95.5 (0.2)
CINIC 94.5 (0.2)
CIFAR-10.1 94.8 (0.3)
CIFAR-10.2 95.3 (0.3)
Table 6: AutoFT outperforms fine-tuning with several different validation sets. We fine-tune on CIFAR-10 and use $100$ examples from either the ID distribution (CIFAR-10) or a non-ID distribution (CIFAR-10-C, CINIC, CIFAR-10.1, CIFAR-10.2) to learn the objective and hyperparameters. Evaluation is on held-out examples from CIFAR-10.1 and CIFAR-10.2.
6.1 Distribution Shifts
Improvements in OOD generalization. We evalute AutoFT on three common robust fine-tuning benchmarks: ImageNet, WILDS-FMoW, and WILDS-iWildCam. This evaluation includes diverse distribution shift conditions including subpopulation and domain shifts. ID and OOD performance reported in Table 4 show that AutoFT consistently outperforms existing methods for robust fine-tuning in OOD metrics. Furthermore, these gains are maintained when ensembling initial and fine-tuned models, suggesting that the benefits from AutoFT are complementary to that of ensembling (Wortsman et al., 2022b, a). We show additional experiments on CIFAR-10 and related distribution shifts in Table 11, which show the same conclusion.
State-of-the-art performance on iWildCam and FMoW. To further assess whether the robustness gains from AutoFT scale to larger foundation models, we evaluate AutoFT with a ViT-L/14@336px model, following the largest scale experiment in Goyal et al. (2022a). Results in Table 2 show that AutoFT improves upon the prior state-of-the-art method, FLYP (Goyal et al., 2022a), by $6.0\%$ in OOD macro-F1. Furthermore, on WILDS-FMoW Table 3, AutoFT outperforms a Transformer-specific parameter freezing method (Kumar et al., 2022b), the prior state-of-the-art, by $1.5\%$ in OOD worst-region accuracy. Additionally, AutoFT outperforms the compute-intensive Model Soups (Wortsman et al., 2022a) on both benchmarks; this method ensembles more than $70$ models fine-tuned with different augmentations and hyperparameters.
6.2 Few-Shot Classification
Few-shot classification serves as an important benchmark in applications where there are small quantities of labeled data available for fine-tuning. Few-shot binary classification is particularly challenging, given the small number of training examples. We evaluate on 4, 16, and 32 shot binary classification tasks from the PatchCamelyon and Rendered-SST2 datasets, following Radford et al. (2021b) and Goyal et al. (2022a). PatchCamelyon contains digital pathology images for the detection of metastatic tissue. Rendered-SST2 focuses on optical character recognition for classifying text sentiment.
Figure 5 shows that AutoFT still improves effective robustness with limited data, outperforming FLYP by $3.7\%$ and full fine-tuning by $3.9\%$ in a challenging 4-shot classification task on Rendered-SST2.
Figure 4: In binary few-shot classification, AutoFT outperforms existing robust fine-tuning methods. AutoFT outperforms FLYP by $3.1\%$ and full fine-tuning by $4.3\%$ in 32-shot classification on PatchCamelyon.
@@ 6.2 Few-Shot Classification
Figure 4: In binary few-shot classification, AutoFT outperforms existing robust fine-tuning methods. AutoFT outperforms FLYP by $3.1\%$ and full fine-tuning by $4.3\%$ in 32-shot classification on PatchCamelyon.
Figure 5: Test performance of AutoFT on WILDS-iWildCam with varying validation set sizes. Increasing size beyond $1000$ shows minimal additional benefits.
6.3 Transfer Learning
We additionally evaluate AutoFT on standard transfer learning tasks, where the evaluation data is drawn from the same distribution as both the training and validation data. We compare AutoFT against other fine-tuning methods on several transfer learning datasets: CalTech-101, Stanford Cars, and Flowers-102. Results in Table 10 demonstrate that AutoFT is competitive with these baselines, but the advantages from learning an objective are less pronounced compared to previous experiments. This suggests that problem-specific learned fine-tuning objectives are most helpful in settings with high underspecification, such as OOD evaluation conditions or few-shot learning, confirming our initial intuition in Section 4.
6.4 Analysis
Transferability of the learned objective. We evaluate models fine-tuned on WILDS-iWildCam with an objective learned on the same WILDS-iWildCam task (AutoFT) or a different task, WILDS-FMoW (AutoFT, transferred objective). We additionally compare with standard fine-tuning (FT) in Table 5. Fine-tuning with an objective learned on a different task (AutoFT, transferred objective) degrades performance, suggesting that the learned objective is not universal, but tailored to the task on which it is optimized.
Choice of dataset for evaluating the learned hyperparameters.
Validation Set ID Test OOD Test
ID Validation 52.3 35.8
OOD Validation 51.0 38.3
Table 7: Using a non-ID validation set improves OOD generalization compared to using an ID validation set. We evaluate AutoFT on WILDS-iWildCam, using either the ID or OOD validation split from the WILDS benchmark to learn the objective and hyperparameters. Using an ID validation set improves ID performance, while using a non-ID validation set improves OOD performance.
In Table 7, we run AutoFT with different validation sets – the official ID validation, OOD validation, and OOD test splits from the WILDS-iWildCam benchmark. Optimizing hyperparameters with respect to an ID validation set improves ID performance, while using a non-ID validation set improves OOD generalization.
Choice of hyperparameter optimization algorithm.
Hyperparameter Optimization ID Test OOD Test
Random 29.8 24.5
Quasi Monte Carlo 48.2 34.7
Bayesian Optimization 49.9 36.5
Tree-Structured Parzen Estimator 51.0 38.3
Table 8: We run AutoFT with various hyperparameter optimization algorithms given the same computational budget. Evaluation is on WILDS-iWildCam. The Tree-Structured Parzen Estimator (TPE) outperforms other methods by a large margin.
We assess the effect of different hyperparameter optimization algorithms on the performance of AutoFT. We run AutoFT with three different hyperparameter optimization methods: random search, quasi Monte Carlo, and Bayesian optimization. Results in Table 8 show that Tree-Structured Parzen Estimator (TPE) outperforms other hyperparameter optimization methods.
Effect of validation set size. We examine the effect of validation set size on performance in Figure 5. Across all validation set sizes, AutoFT yields improvements in OOD performance compared to the leading fine-tuning baseline (FLYP: $35.6\%$). Given a fixed number of inner steps and hyperparameter evaluations, there’s an optimal range for validation set size, beyond which additional benefits are minimal.
7 Conclusion
We introduce AutoFT, a novel, data-driven approach for robust fine-tuning that learns the objective and hyperparameters. With a small amount of data from one OOD distribution, which is often readily available, AutoFT outperforms prior fine-tuning methods. Notably, AutoFT attains state-of-the-art performance on two WILDS benchmarks.
Limitations. AutoFT is a general recipe for robust adaptation, and our specific implementation is only one instantiation. We expect that future work can improve upon AutoFT by investigating other loss parameterizations and meta-optimization techniques within the AutoFT framework. While our experiments show strong results in image classification, we have not yet evaluated AutoFT in other problem settings, such as natural language processing (NLP). We hope that our work will inspire future work on data-driven approaches for robust adaptation.
Acknowledgements
We thank Kyle Hsu, Lukas Haas, and members of the IRIS lab for helpful feedback and discussions. We also thank Sachin Goyal for help with ImageNet experiments. This work was supported by KFAS and NSF. A.R. is supported by Schmidt Futures, Google, Apple, and Open Philanthropy. C.F. is supported by ONR grant N00014-21-1-2685 and the NSF CAREER award.
Appendix A Intuition and Didactic Experiment for AutoFT
To illustrate how hyperparameter optimization can lead to more effective adaptation of a partially useful prior, we consider a simple experiment. We focus on synthetic datasets of vectors drawn from zero-mean Gaussian distributions with different variances.
We assume that the prior unit variance distribution is helpful but that the data exhibits deviation from this prior along a few dimensions. The prior distribution is a unit Gaussian, i.e., it has variance $[{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},\ldots]$. Training data $D_{\text{tr}}$ is drawn from a distribution with variance $[{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},\ldots]$, while the validation data $D_{\text{val}}$ is drawn from a distribution with variance $[{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},\ldots]$. We fit the parameters of a Gaussian distribution $q_{\theta}$ to the training data using a variational Bayes objective by minimizing the following loss:
We consider “fine-tuning” as starting from the prior distribution. In this toy model, dimension-wise weight hyperparameters $w_{i}$ dictate the degree of influence of the prior versus the training data on each dimension, analogous to how the fine-grained hyperparameters in AutoFT balance the foundation model prior and training data along different aspects of fine-tuning. As in AutoFT, we use the TPE algorithm to optimize the weight hyperparameters $w_{i}$ to maximize the log-likelihood of the validation data $D_{\text{val}}$.
Appendix B Additional Results
B.1 ImageNet Distribution Shifts
B.1.1 Main Results
We provide detailed results for each ImageNet-derived distribution shift in Table 9. AutoFT outperforms all baselines on 9 out of 10 OOD datasets. AutoFT improves OOD performance by $3.0\%$ compared to the leading baseline, FLYP. We observe a similar effect with all other points of comparison, suggesting that perhaps these prior methods for robust fine-tuning “overfit” to the ID distribution in a way that AutoFT does not.
Without Ensembling With Ensembling
Methods ID Im-V2 Im-R Im-A Sketch ObjectNet Avg. OOD ID Im-V2 Im-R Im-A Sketch ObjectNet Avg. OOD
Zeroshot 68.3 61.9 77.7 50.0 48.3 55.4 58.7 68.3 61.9 77.7 50.0 48.3 55.4 58.7
LP 79.9 69.8 70.8 46.4 46.9 52.1 57.2 80.0 70.3 72.4 47.8 48.1 52.8 58.3
FT 81.3 71.2 66.1 37.8 46.1 53.3 54.9 82.5 72.8 74.9 48.1 51.9 59.0 61.3
L2-SP 81.7 71.8 70.0 42.5 48.5 56.2 57.8 82.2 72.9 75.1 48.6 51.4 58.9 61.4
LP-FT 81.7 72.1 73.5 47.6 50.3 58.2 60.3 82.1 72.8 75.3 50.1 51.7 59.2 61.8
FLYP 82.6 73.0 71.4 48.1 49.6 58.7 60.2 82.9 73.5 76.0 53.0 52.3 60.8 63.1
AutoFT 81.8 73.1 72.4 48.8 49.8 63.5 61.5 82.4 73.6 76.4 53.1 52.6 65.6 64.3
Table 9: Detailed ImageNet results. AutoFT consistently outperforms all baselines on 9 out of 10 ImageNet distribution shifts.
B.1.2 Few-Shot Classification
Figure 7 plots effective robustness at varying interpolation coefficients on five ImageNet distribution shifts for 4, 16, and 32 shot classification. In all three settings, AutoFT results in better OOD generalization and avoids overfitting to the small fine-tuning dataset. For example, in 32-shot ImageNet, AutoFT improves average OOD accuracy by $2.2\%$ compared to the leading baseline, FLYP.
Figure 7: AutoFT improves OOD generalization in few-shot learning. AutoFT results in gains in effective robustness over baselines, indicated by vertical distance. AutoFT results in better OOD performance at a given ID performance threshold across varying degrees of weight interpolation between the fine-tuned and zero-shot models.